Parramore Corp has $10 million of sales, $1 million of inventories, $4 million of receivables, and $2 million of payables. Its cost of goods sold is 85% of sales, and it finances working capital with bank loans at an 8% rate. Assume 365 days in year for your calculations. Do not round intermediate steps.

  1. What is Parramore's cash conversion cycle (CCC)? Do not round intermediate calculations. Round your answer to two decimal places.
      days

  2. If Parramore could lower its inventories and receivables by 11% each and increase its payables by 11%, all without affecting sales or cost of goods sold, what would be the new CCC? Do not round intermediate calculations. Round your answer to two decimal places.
      days

  3. How much cash would be freed up, if Parramore could lower its inventories and receivables by 11% each and increase its payables by 11%, all without affecting sales or cost of goods sold? Do not round intermediate calculations. Round your answer to the nearest cent. Write out your answer completely. For Example, 13.2 million should be entered as 13,200,000.
    $

  4. By how much would pretax profits change, if Parramore could lower its inventories and receivables by 11% each and increase its payables by 11%, all without affecting sales or cost of goods sold? Do not round intermediate calculations. Round your answer to the nearest cent. Write out your answer completely. For Example, 13.2 million should be entered as 13,200,000.
    $
Answer #1

Sales = $10 million

Inventory = $ 1 million

Receivable = $ 4 million

Payables = $ 2 million

COGS = 85% of Sales

Loan rate = 8%

1)

Cash conversion cycle (CCC) = Days inventory outstanding (DIO) + Days sales outstanding (DSO) - Days payable outstanding (DPO)

Days inventory outstanding (DIO) = (Inventory / COGS) *365 = (1 / (85%*10))*365 = 42.941176

Days sales outstanding (DSO) =  (Accounts receivable / Total credit sales) * 365 = (4/10)*365 = 146

Days payable outstanding (DPO) = (Accounts payable / COGS) * 365 = (2/(85%*10)*365 = 85.88235

Cash conversion cycle (CCC) = 42.941176 + 146 - 85.88235 = 103.0588235

CCC is 104 days

2) Inventories reduced by 11%, Inventory = $1 million *(1-11%) = $ 0.89 million

Receivables reduced by 11%, Receivable = $ 4 million *(1-11%) = $ 3.56 million

Payable increased by 11%, Payable = $2 million*(1+11%) = $ 2.2 million

Cash conversion cycle (CCC) = Days inventory outstanding (DIO) + Days sales outstanding (DSO) - Days payable outstanding (DPO)

Days inventory outstanding (DIO) = (Inventory / COGS) *365 = (0.89 / (85%*10))*365 = 38.21764

Days sales outstanding (DSO) =  (Accounts receivable / Total credit sales) * 365 = (3.56/10)*365 = 129.94

Days payable outstanding (DPO) = (Accounts payable / COGS) * 365 = (2.2/(85%*10)*365 = 95.3294

Cash conversion cycle (CCC) = 38.2176 + 129.94- 95.3294 = 72.8282

CCC is 73 days

3) Cash Freed up = Cash freed up for inventory + Cash freed up for receivable + Cash freed up for payable

Cash freed up for inventory = Earlier Inventory - Inventory after reduction = 1-0.89 = 0.11 miliion = $110,000

Cash freed up for receivable = Earlier Receivable - Receivable after reduction = 4 - 3.56 = 0.44 million = $440,000

Cash freed up for payable = Earlier payable - payable after reduction = 2 - 2.22 = - 0.22 million = - $220,000

Cash Freed up = $110,000 + $440,000 - $220,000 = $330,000

4) Pretax profit from freed up capital = $330,000 * 8% = $26,400